A University wishes to hire 4 faculty members for its newly established department. They have to select from a group of 8 Male and 7 Female applicants, all of whom are equally qualified to fill in the positions. If University selects 4 at random, what is the probability that:
a) All four will be Men
b) All four will be Women
c) Atleast one will be a Women
d) Atleast one will be a Men
1
Expert's answer
2020-06-11T20:57:44-0400

Total No of faculty "= 8 M + 7 F = 15"


 If University selects 4 at random,


a) All four will be Men "= \frac{_8C_4}{_{15}C_4}= \frac{2}{39}"


b) All four will be Women "= \frac{_7C_4}{_{15}C_4}= \frac{1}{39}"


c) At least one will be a Women = 1 - Probability of all men "= 1-\frac{_8C_4}{_{15}C_4}= \frac{37}{39}"

d) At least one will be a Men = 1 - Probability of all women

"= 1-\frac{_7C_4}{_{15}C_4}= \frac{38}{39}"